OpenStudy (anonymous):

x+y=-2 x+y=3 the two equation notice the left sides of the equation are exactly alike, but the right are different. this mean the to lines are parallel and they dont intersect. there fore there is no solution in this system. then i have to solve using the substitution we will do that first . solve for x????? i dont know CAn you guys help me? Thanks :)

OpenStudy (anonymous):

You are correct, these are two parallel lines. One of them is 5 units above the other. What are you trying to find?

OpenStudy (anonymous):

Sovle for the x

OpenStudy (anonymous):

Well, in the first equation x = -2-y And in the second: x=3-y

OpenStudy (anonymous):

yeah, keep going please

OpenStudy (anonymous):

That's it: -2-y=3-y -2=3 Which is impossible.

OpenStudy (anonymous):

So what should i do? I still kindas not really get it

OpenStudy (anonymous):

Based on your current question there is no answer. Is this something you made up or can you type in the exact question?

OpenStudy (anonymous):

Oh wait never mind i get it, my bad sorry. Yeah you was right, thanks